# WaterSeeker: Efficient Detection of Watermarked Segments in Large Documents

### v2

@@ 8 Conclusion
In this work, we defined a new scenario for detecting watermarked segments in large documents and established its evaluation metrics. We identified the limitations of traditional full-text detection methods in this context and proposed a “first locate, then detect” watermark detection algorithm that utilizes a coarse-to-fine strategy. We validated the detection performance and time complexity of our algorithm through a series of analyses and experiments, demonstrating its ability to effectively balance both aspects. For future improvements, we recommend exploring more advanced locating methods based on this concept, which may yield better detection results.
Appendix A Pseudocode of Detection Baselines
@@
Appendix B Detailed Proof for Aar
B.1 Expected p-value of watermarked text
@@
$E[\text{p-value}]=E[\text{GammaTransform}(S,W,\text{loc}=0,\text{scale}=1)]$ (9)
@@
$E[\text{GammaTransform}(S,W,0,1)]\geq\text{GammaTransform}(E[S],W,0,1)$ (10)
Therefore, we can analyze $\text{GammaTransform}(E[S],W,0,1)$ to obtain a lower bound for the expectation of p-value. Define function $f(W)=\text{GammaTransform}(E[S],W,0,1)$:
  • When $W\leq L$: $f(W)=\text{GammaTransform}(W\mu_{1},W,0,1)$
  • When $W>L$: $f(W)=\text{GammaTransform}(L\mu_{1}+(W-L)\mu_{0},W,0,1)$
@@
B.2 Calculation of p-threshold to control false positive rate
@@
Appendix C Detail of Threshold Selection Within the Specified Window
A key role of threshold selection is to control the false positive rate. In this context, the task involves detecting watermark fragments within long texts, which requires traversing extensive content and can lead to an accumulation of false positives. Therefore, managing the false positive rate within the detection window is crucial in this scenario. In the experiment, we set the targeted false positive rate $\alpha$ within the detection window to $10^{-6}$.
C.1 Rationale for Setting $\alpha$ to $10^{-6}$
@@
WaterSeeker, WinMax, and FLSW all involve employing sliding windows for text traversal and conduct full-text detection within each window. As these windows overlap, they cannot be treated as independent, making it challenging to derive a theoretical upper bound for the document-level FPR from the targeted FPR within each window. Given this, we utilize large-scale data simulation to demonstrate that, with a targeted false positive rate of $10^{-6}$ within each window, our proposed method WaterSeeker maintains an acceptable false positive rate.
@@
We then apply WaterSeeker to detect watermarked segments within these samples, setting the targeted false positive rate within the detection window to $10^{-6}$. The large-scale simulation results in Table 5 demonstrate that WaterSeeker maintains a false positive rate of approximately 0.005, which is considered acceptable. For scenarios requiring more stringent FPR control, the targeted false positive rate can be adjusted downward. However, this inevitably compromises the detection rate, highlighting a key challenge in watermarked segment detection within large documents.
C.2 Setting the Threshold to Achieve a Target False Positive Rate $\alpha$
@@
For Aar, as analyzed in Appendix B, $p^{*}=\alpha$, therefore set $p^{*}=10^{-6}$ for all lengths.
Appendix D Sample Distribution in the Main Experiment

### v3

@@ 7 Conclusion
This work introduces a new scenario for detecting watermarked segments in large documents and establishes corresponding evaluation metrics. We identified the limitations of full-text detection methods in this context and proposed a “first locate, then detect" watermark detection algorithm that utilizes a coarse-to-fine strategy. We validated the detection performance and time complexity of our algorithm through a series of analyses and experiments, demonstrating its ability to effectively balance both aspects. Future research could explore more advanced locating methods based on this concept to potentially yield improved detection results.
Limitations
While our method has demonstrated effectiveness in detecting watermarked segments within large documents, there are still some limitations that need to be addressed in future work. First, from an evaluation perspective, due to resource constraints, we only conducted experiments on Llama-2-7B and Mistral-7B models. The effectiveness of our method on larger and more powerful models remains to be further verified. Second, WaterSeeker’s performance may decrease with very short or weak watermarks. Enhancing the sensitivity of WaterSeeker to detect shorter and weaker watermarks is an area for future improvement, which may involve refining the anomaly extraction algorithms or incorporating additional contextual analysis. Lastly, parameter selection, such as threshold settings, is crucial and can be challenging in different environments. Currently, parameters are manually tuned based on observations, so developing adaptive tuning methods, potentially using machine learning, would enhance WaterSeeker’s robustness and practicality across diverse scenarios.
Appendix A Details of Representative Watermarking Algorithms
A.1 KGW
Watermarking. In watermarked text generation, the process for the $t$-th token begins by hashing preceding tokens with a secret key, creating a red-green vocabulary partition where green tokens comprise a fraction $\gamma$. Green token logits are then incrementally increased by $\delta$, which can be expressed as follows:
$l^{\prime}_{t}(y)=\begin{cases}l_{t}(y),&y\in R_{t}\\ l_{t}(y)+\delta,&y\in G_{t}\end{cases}$ (1)
This subtle modification results in watermarked text exhibiting a higher frequency of green tokens compared to non-watermarked text.
Detection. Detecting a KGW watermark entails computing red-green partitions for each position using preceding tokens and the hash function, then calculating the green token proportion using the z-score:
$z=\frac{|s|_{G}-\gamma N}{\sqrt{\gamma(1-\gamma)N}}$ (2)
, where $|s|_{G}$ represents the total count of green tokens in the whole text of length $N$.
A.2 Aar
Watermarking. When generating the $t$-th token, it first involves hashing the preceding tokens using a secret key to obtain a pseudo vector $u_{t}\sim\text{Uniform}([0,1])^{|V|}$. The $t$-th token is determined by
$\mathop{\arg\max}\limits_{y}\,u_{t}(y)^{1/p_{t}(y)},$ (3)
where $p$ is the probability vector produced by LLM at the $t$-th step. Let’s perform equivalent transformations on it:
$\bf{y}$ $=\mathop{\arg\max}\limits_{y}\,u_{t}(y)^{1/p_{t}(y)}$ (4)
$=\mathop{\arg\max}\limits_{y}\,\frac{1}{p_{t}(y)}\log u_{t}(y)$ (5)
$=\mathop{\arg\min}\limits_{y}\,\frac{1}{p_{t}(y)}\log\frac{1}{u_{t}(y)}$ (6)
$=\mathop{\arg\min}\limits_{y}\,\log\frac{1}{p_{t}(y)}+\log\log\frac{1}{u_{t}(y)}$ (7)
$=\mathop{\arg\max}\limits_{y}\,\log p_{t}(y)-\log\log\frac{1}{u_{t}(y)}$ (8)
Given that the probabilistic output $p_{t}$ of an LLM is derived from the logits $l_{t}$ through a softmax transformation, and when we additionally consider the sampling temperature $T$, Equation 8 becomes equivalent to:
$\mathop{\arg\max}\limits_{y}\,\cfrac{l_{t}(y)}{T}+G_{t}(y),$ (9)
where $l_{t}$ is the logits produced by the LLM, and $G_{t}$ is the Gumbel noise: $G_{t}(y)\sim Gumbel(0,1)$. The $Gumbel(0,1)$ distribution is defined as follows: if $u\sim\text{Uniform}(0,1)$, then $-\log(-\log(u))$ follows a $Gumbel(0,1)$ distribution.
It is evident that the temperature $T$ can be utilized to exert control over the watermark strength. As the value of $T$ increases, the influence of Gumbel noise on the sampling process becomes more pronounced, consequently resulting in a stronger watermark.
Detection. Detecting an Aar watermark involves calculating the correlation value between the pseudo vector $u_{t}$ and the corresponding token $y_{t}$ in the text to be examined. The correlation value can be expressed as:
$\log\cfrac{1}{1-u_{t}(y_{t})}.$ (10)
For the entire text, the statistic value can be expressed as:
[…]
where $\Gamma$ is the Gamma Transformation function that converts the sum of correlation values to a p-value.
Appendix B Pseudocode of WaterSeeker and Detection Baselines
Pseudocode of WaterSeeker, WinMax and FLSW could be found in Algorithm 3, 1 and 2, respectively.
@@ Appendix B Pseudocode of WaterSeeker and Detection Baselines
Algorithm 3 WaterSeeker Algorithm
2:     Compute score list using sliding window
3:     Calculate $s_{\text{mean}}$, $s_{\text{top-k-mean}}$, and $s_{\text{std}}$
4:     Detect anomalies using threshold:
6:     Connect nearby outliers and filter short segments
7:     return filteredSegments
8:end procedure
10:     hasWatermark $\leftarrow$ False
11:     indices $\leftarrow$ [ ]
12:     for $(s^{\prime},e^{\prime})$ in suspiciousRegions do
13:         maxScore = $\infty$- ∞, maxIndice = ()
15:              if DetectWatermark(tokens[$s:e$])$>$ maxScore then
16:                  maxScore = DetectWatermark(tokens[$s:e$])
17:                  maxIndice = $(s,e)$
18:              end if
19:         end for
20:         if maxScore $>$ threshold then
21:              hasWatermark $\leftarrow$ True
22:              indices.append(maxIndice)
23:         end if
24:     end for
25:     return hasWatermark, indices
26:end procedure
30:     return hasWatermark, indices
31:end procedure
Appendix C Detailed Proof for Aar
C.1 Expected p-value of watermarked text
@@
$E[\text{p-value}]=E[\Gamma(S,W,\text{loc}=0,\text{scale}=1)]$ (13)
@@
$E[\Gamma(S,W,0,1)]\geq\Gamma(E[S],W,0,1)$ (14)
Therefore, we can analyze $\Gamma(E[S],W,0,1)$ to obtain a lower bound for the expectation of p-value. Define function $f(W)=\Gamma(E[S],W,0,1)$:
  • When $W\leq L$: $f(W)=\Gamma(W\mu_{1},W,0,1)$
  • When $W>L$: $f(W)=\Gamma(L\mu_{1}+(W-L)\mu_{0},W,0,1)$
@@
C.2 Calculation of p-threshold to control false positive rate
@@
Appendix D Detail of Threshold Selection Within the Specified Window
A key role of threshold selection is to control the false positive rate. In this context, the task involves detecting watermark fragments within long texts, which requires traversing extensive content and can lead to an accumulation of false positives. Therefore, managing the false positive rate within the detection window is crucial in this scenario. In the experiment, we set the target false positive rate $\alpha$ within the detection window to $10^{-6}$.
D.1 Rationale for setting $\alpha$ to $10^{-6}$
@@
WaterSeeker, WinMax, and FLSW all involve employing sliding windows for text traversal and conduct full-text detection within each window. As these windows overlap, they cannot be treated as independent, making it challenging to derive a theoretical upper bound for the document-level FPR from the target FPR within each window. Given this, we utilize large-scale data simulation to demonstrate that, with a target false positive rate of $10^{-6}$ within each window, our proposed method WaterSeeker maintains an acceptable false positive rate.
@@
We then apply WaterSeeker to detect watermarked segments within these samples, setting the target false positive rate within the detection window to $10^{-6}$. The large-scale simulation results in Table 5 demonstrate that WaterSeeker maintains a false positive rate of approximately 0.005, which is considered acceptable. For scenarios requiring more stringent FPR control, the target false positive rate can be adjusted downward. However, this inevitably compromises the detection rate, highlighting a key challenge in watermarked segment detection within large documents.
D.2 Setting the threshold to achieve a target false positive rate $\alpha$
@@
For Aar, as analyzed in Appendix C, $p^{*}=\alpha$, therefore set $p^{*}=10^{-6}$ for all lengths.
Appendix E Impact of Window Size on Watermark Intensity Calculation
(a) $W=1$
(b) $W=10$
(c) $W=30$
(d) $W=50$
Figure 4: Case study: Impact of varying window sizes on watermark intensity calculation in the KGW algorithm.
(a) $W=1$
(b) $W=10$
(c) $W=30$
(d) $W=50$
Figure 5: Case study: Impact of varying window sizes on watermark intensity calculation in the Aar algorithm.
The first step in WaterSeeker is score list computation. In this step, selecting an appropriate window size $W$ for calculating mean scores is crucial. A small $W$ introduces excessive noise, while a large $W$ reduces granularity and increases computational time due to the need to examine $W^{2}$ windows during local traversal. Therefore, we aim to determine an appropriate window size that is relatively small while still providing a sufficiently smooth representation of watermark intensity throughout the text.
We present a case study comparing watermark intensity calculations using window sizes $W=1,10,30,$ and $50$. The analysis encompasses the ground truth segment and 500 tokens on either side. Figures 4 and 5 illustrate the results for the KGW and Aar algorithms, respectively. The intensity curves reveal that small window sizes, particularly $W\leq 10$, introduce significant fluctuations. While $W=30$ exhibits reduced noise, it still presents instabilities, as shown in Figure 5(c) (the ground truth segment part). Overall, $W=50$ demonstrates the least noise. Consequently, we adopt $W=50$ for our main experiments.
Appendix F Sample Distribution in the Main Experiment
@@
Appendix G Further Experimental Results under Different Mixing Ratios
In main experiment, 100-400 watermarked tokens are inserted into non-watermarked text of 10,000 tokens, resulting in watermarked text mixing ratios of 1% to 4%. This section expands the evaluation by testing the performance of various methods across a broader range of mixing ratios, achieved by adjusting the length of the non-watermarked text. We deliberately avoid adjusting the mixing ratio by inserting multiple watermark segments into the non-watermarked text, as the WinMax method can only return the highest-scoring segment and is not designed for multi-segment insertion scenarios. Additionally, we do not use fixed-length watermark segments because in real-world detection scenarios, the length of the watermark segment is typically unknown. Moreover, if the watermark length were fixed and known, a simple Fixed-length Sliding Window detection method (with the window size equal to the watermark segment length) would suffice to solve the problem.
We provided further experimental results under three different settings:
  1. 1.
    100-400 watermarked tokens are inserted into 5,000 non-watermarked tokens, with a mixing ratio of 2% (100 / 5100) -7.4% (400 / 5400). The results are demonstrated in Table 6.
  2. 2.
    100-400 watermarked tokens are inserted into 2,000 non-watermarked tokens, with a mixing ratio of 4.8% (100 / 2100) -16.7% (400 / 2400). The results are demonstrated in Table 7.
  3. 3.
    100-400 watermarked tokens are inserted into 500 non-watermarked tokens, with a mixing ratio of 16.7% (100 / 600) -44.4% (400 / 900). The results are demonstrated in Table 8.
Table 6: Comparison of WaterSeeker with other baselines under the mixing ratio of 2%-7.4%.
Model Method KGW Aar
FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
Llama-2-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.001 0.000 1.000 [0.957] 0.000 [0.083] 0.002
WinMax 0.003 0.317 0.810 0.621 0.010 0.380 0.761 0.565
FLSW-100 0.000 0.520 0.649 0.413 0.003 0.523 0.644 0.375
FLSW-200 0.000 0.457 0.704 0.404 0.000 0.560 0.611 0.350
FLSW-300 0.003 0.663 0.502 0.297 0.000 0.767 0.378 0.283
FLSW-400 0.003 0.867 0.235 0.230 0.000 0.920 0.148 0.205
WaterSeeker(Ours) 0.003 0.357 0.781 0.570 0.010 0.407 0.740 0.536
Mistral-7b Full-text Detection 0.000 1.000 [0.990] 0.000 [0.020] 0.000 0.000 1.000 [0.937] 0.000 [0.119] 0.004
WinMax 0.003 0.227 0.871 0.700 0.010 0.370 0.768 0.570
FLSW-100 0.000 0.503 0.664 0.443 0.000 0.543 0.627 0.364
FLSW-200 0.000 0.350 0.788 0.497 0.003 0.540 0.629 0.362
FLSW-300 0.000 0.527 0.643 0.392 0.003 0.763 0.382 0.269
FLSW-400 0.000 0.783 0.357 0.281 0.003 0.910 0.165 0.203
WaterSeeker(Ours) 0.000 0.243 0.861 0.676 0.010 0.383 0.758 0.548
Table 7: Comparison of WaterSeeker with other baselines under the mixing ratio of 4.8%-16.7%.
Model Method KGW Aar
FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
Llama-2-7b Full-text Detection 0.000 1.000 [0.967] 0.000 [0.065] 0.005 0.000 1.000 [0.873] 0.000 [0.225] 0.018
WinMax 0.003 0.257 0.851 0.669 0.003 0.413 0.738 0.533
FLSW-100 0.000 0.473 0.690 0.440 0.003 0.570 0.600 0.350
FLSW-200 0.000 0.420 0.734 0.446 0.000 0.607 0.565 0.327
FLSW-300 0.003 0.630 0.539 0.339 0.000 0.777 0.365 0.257
FLSW-400 0.003 0.850 0.260 0.246 0.000 0.883 0.209 0.200
WaterSeeker(Ours) 0.003 0.287 0.831 0.642 0.003 0.443 0.714 0.501
Mistral-7b Full-text Detection 0.000 1.000 [0.980] 0.000 [0.039] 0.003 0.000 1.000 [0.870] 0.000 [0.230] 0.017
WinMax 0.000 0.220 0.876 0.710 0.003 0.353 0.783 0.590
FLSW-100 0.000 0.470 0.693 0.447 0.000 0.487 0.678 0.415
FLSW-200 0.000 0.360 0.780 0.497 0.003 0.500 0.665 0.371
FLSW-300 0.000 0.533 0.636 0.386 0.003 0.760 0.386 0.285
FLSW-400 0.000 0.747 0.404 0.288 0.003 0.887 0.203 0.212
WaterSeeker(Ours) 0.000 0.227 0.872 0.701 0.003 0.363 0.776 0.574
Table 8: Comparison of WaterSeeker with other baselines under the mixing ratio of 16.7%-44.4%.
Model Method KGW Aar
FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$ FPR $\downarrow$ FNR $\downarrow$ F1 $\uparrow$ IoU $\uparrow$
Llama-2-7b Full-text Detection 0.000 1.000 [0.787] 0.000 [0.352] 0.081 0.000 1.000 [0.693] 0.000 [0.469] 0.109
WinMax 0.000 0.247 0.859 0.685 0.000 0.373 0.770 0.568
FLSW-100 0.000 0.470 0.693 0.445 0.000 0.557 0.614 0.380
FLSW-200 0.000 0.383 0.763 0.461 0.000 0.553 0.618 0.340
FLSW-300 0.000 0.553 0.618 0.360 0.000 0.690 0.473 0.298
FLSW-400 0.000 0.740 0.413 0.265 0.000 0.837 0.281 0.231
WaterSeeker(Ours) 0.000 0.277 0.839 0.647 0.000 0.410 0.742 0.522
Mistral-7b Full-text Detection 0.000 1.000 [0.847] 0.000 [0.266] 0.061 0.000 1.000 [0.697] 0.000 [0.465] 0.112
WinMax 0.000 0.270 0.844 0.652 0.000 0.350 0.788 0.588
FLSW-100 0.000 0.510 0.658 0.432 0.000 0.530 0.639 0.393
FLSW-200 0.000 0.407 0.745 0.442 0.000 0.510 0.658 0.369
FLSW-300 0.000 0.570 0.601 0.352 0.000 0.690 0.473 0.294
FLSW-400 0.000 0.760 0.387 0.243 0.000 0.827 0.295 0.246
WaterSeeker(Ours) 0.000 0.303 0.821 0.612 0.000 0.407 0.745 0.543
Experimental results demonstrate that WaterSeeker exhibits high robustness across various mixing ratios, showing consistent performance trends across different settings. Specifically, WaterSeeker significantly outperforms full-text detection and FLSW methods in both detection accuracy and localization effectiveness, while achieving comparable performance to the upper bound method, WinMax.
Appendix H Further Application Scenario